# Emergent time scales of epistasis in protein evolution

### 14 Mar 2024 v1

@@ Abstract
Figure 3: Contingency and entrenchment from in silico evolution for the $\beta$-lactamase family. (A) Scatter plot over sites of distinct notions of mutability. We first select 5000 mutations labeled by $aib$ (i.e. site $i$, aminoacid $a\to b$) along an evolutionary trajectory of the $\beta$-lactamase family. On the horizontal axis is reported for each $aib$ the CDE of site $i$ in the context corresponding to the time at which the sites mutates, i.e. $t=t_{\rm MUT}$. On the vertical axis are reported both the mean CDE on all natural proteins in the family (blue) and the CIE (red) of the site $i$. (B) Contingency: average value of the fitness gain $-\Delta\Delta E$ between the original mutation $aib$ at time $t_{\rm MUT}$, and the mutational effect of introducing aminoacid $b$ in site $i$ over a time window of 5000 MCMC steps before $t_{\rm MUT}$. The faded area is the standard deviation. (C) Entrenchment: average value of the fitness gain $-\Delta\Delta E$ between the original mutation $aib$ at time $t_{\rm MUT}$, and the mutational effect of reverting to aminoacid $a$ over a time window of 5000 MCMC steps after $t_{\rm MUT}$. The faded area is the standard deviation.
@@
To further test this idea, we first consider the negative variation of energy, $-\Delta E(t=t_{\rm MUT})$, associated to a given mutation at the time $t_{\rm MUT}$ at which it is first observed along an evolutionary trajectory. Fitness is negatively correlated to energy, such that a positive $-\Delta E$ indicates a beneficial mutation. We find that along the evolutionary trajectories, the distribution of $-\Delta E(t=t_{\rm MUT})$ is centered around zero, with symmetric tails at positive and negative values. This means that observed mutations are close to neutral, as opposed to a strongly deleterious bias of random mutations (remember that our evolutionary model combines mutation and selection in its elementary MCMC step). This is consistent with the DCA energy remaining on average constant along trajectories, but also with the empirical observation that natural aminoacid polymorphisms tend to be neutral [17] when scored by a DCA model. To better quantify this effect and to reduce statistical noise, we consider
${\rm CDE}^{\rm mut}_{i}\equiv{\rm CDE}_{i}[A_{\setminus i}(t=t_{\rm MUT})]\ ,$ (4)
i.e. the CDE at a given site $i$, calculated at the mutation time and using the background in which the mutation occurs. For the same site $i$, we also consider ${\rm CIE}_{i}$ as given by (2), and the average CDE over the natural MSA
$\overline{{\rm CDE}}^{\rm nat}_{i}\equiv\frac{1}{|\text{MSA}|}\sum_{A\in\text{MSA}}{\rm CDE}_{i}[A]\ .$ (5)
Fig. 3A shows a scatter plot of these quantities over mutations happening along the trajectory. Remarkably, we observe that, at the time when they mutate, sites have an anomalously high mutability, very close to the overall context-independent mutability but much higher than the average mutability over natural sequences, i.e. ${\rm CDE}^{\rm mut}_{i}\approx{\rm CIE}_{i}>\overline{{\rm CDE}}^{\rm nat}_{i}$. This confirms and quantifies the intuitive idea that, due to purifying selection, when a mutation is introduced, the corresponding site is weakly constrained by the context, and the observed mutations are typically neutral.
@@
The mutational effects are calculated for site $i$, but considering the different backgrounds present at $t_{\rm MUT}$ and $t+t_{\rm MUT}$. Once again, the sign is chosen in such a way that a positive $-\Delta\Delta E$ indicates a gain of fitness. For negative $t$, i.e. times prior to the actual occurrence of the mutation $a\to b$, we test the insertion of aminoacid $b$ into the past sequence $A(t+t_{\rm MUT})$. Fig. 3B shows that the average $-\Delta\Delta E$ is negative when $t<0$ indicating a fitness loss, i.e. a mutation is typically deleterious at previous times, leading to contingency. For positive $t$, i.e. times posterior to the actual occurrence of the mutation $a\to b$, we revert the aminoacid present in $A(t+t_{\rm MUT})$ to $a$. Fig. 3C shows that this reversion is typically deleterious at subsequent times, leading to entrenchment.
@@ B.2 Validation of the model for the DBD family
Figure 10: Contingency and entrenchment from in silico evolution for the DBD family. (A) Scatter plot over sites of distinct notions of mutability. We first select 5000 mutations labeled by $aib$ (i.e. site $i$, aminoacid $a\to b$) along an evolutionary trajectory of the DBD family. On the horizontal axis is reported for each $aib$ the CDE of site $i$ in the context corresponding to the time at which the sites mutates, i.e. $t=t_{\rm MUT}$. On the vertical axis are reported both the mean CDE on all natural proteins in the family (blue) and the CIE (red) of the site $i$. (B) Contingency: average value of the fitness gain $-\Delta\Delta E$ between the original mutation $aib$ at time $t_{\rm MUT}$, and the mutational effect of introducing aminoacid $b$ in site $i$ over a time window of 5000 MCMC steps before $t_{\rm MUT}$. The faded area is the standard deviation. (C) Entrenchment: average value of the fitness gain $-\Delta\Delta E$ between the original mutation $aib$ at time $t_{\rm MUT}$, and the mutational effect of reverting to aminoacid $a$ over a time window of 5000 MCMC steps after $t_{\rm MUT}$. The faded area is the standard deviation.

### 25 Mar 2024 v2

@@ Abstract
Figure 3: Contingency and entrenchment from in silico evolution for the $\beta$-lactamase family. We select 5000 mutations labeled by $aib$ (i.e. site $i$, aminoacid $a\to b$) along an evolutionary trajectory of the $\beta$-lactamase family. (A) Contingency: average value of the fitness gain $-\Delta\Delta E$ between the original mutation $aib$ at time $t_{\rm MUT}$, and the mutational effect of introducing aminoacid $b$ in site $i$ over a time window of 5000 MCMC steps before $t_{\rm MUT}$. The faded area is the standard deviation. (B) Entrenchment: average value of the fitness gain $-\Delta\Delta E$ between the original mutation $aib$ at time $t_{\rm MUT}$, and the mutational effect of reverting to aminoacid $a$ over a time window of 5000 MCMC steps after $t_{\rm MUT}$. The faded area is the standard deviation.
@@
To further test this idea, we first consider the negative variation of energy, $-\Delta E(t=t_{\rm MUT})$, associated to a given mutation at the time $t_{\rm MUT}$ at which it is first observed along an evolutionary trajectory. Fitness is negatively correlated to energy, such that a positive $-\Delta E$ indicates a beneficial mutation. We find that along the evolutionary trajectories, the distribution of $-\Delta E(t=t_{\rm MUT})$ is centered around zero, with symmetric tails at positive and negative values. This means that observed mutations are close to neutral, as opposed to a strongly deleterious bias of random mutations (remember that our evolutionary model combines mutation and selection in its elementary MCMC step). This is consistent with the DCA energy remaining on average constant along trajectories, but also with the empirical observation that natural aminoacid polymorphisms tend to be neutral [17] when scored by a DCA model.
@@
The mutational effects are calculated for site $i$, but considering the different backgrounds present at $t_{\rm MUT}$ and $t+t_{\rm MUT}$. Once again, the sign is chosen in such a way that a positive $-\Delta\Delta E$ indicates a gain of fitness. For negative $t$, i.e. times prior to the actual occurrence of the mutation $a\to b$, we test the insertion of aminoacid $b$ into the past sequence $A(t+t_{\rm MUT})$. Fig. 3A shows that the average $-\Delta\Delta E$ is negative when $t<0$ indicating a fitness loss, i.e. a mutation is typically deleterious at previous times, leading to contingency. For positive $t$, i.e. times posterior to the actual occurrence of the mutation $a\to b$, we revert the aminoacid present in $A(t+t_{\rm MUT})$ to $a$. Fig. 3B shows that this reversion is typically deleterious at subsequent times, leading to entrenchment.
@@ B.2 Validation of the model for the DBD family
Figure 10: Contingency and entrenchment from in silico evolution for the DBD family. We select 5000 mutations labeled by $aib$ (i.e. site $i$, aminoacid $a\to b$) along an evolutionary trajectory of the DBD family. (A) Contingency: average value of the fitness gain $-\Delta\Delta E$ between the original mutation $aib$ at time $t_{\rm MUT}$, and the mutational effect of introducing aminoacid $b$ in site $i$ over a time window of 5000 MCMC steps before $t_{\rm MUT}$. The faded area is the standard deviation. (B) Entrenchment: average value of the fitness gain $-\Delta\Delta E$ between the original mutation $aib$ at time $t_{\rm MUT}$, and the mutational effect of reverting to aminoacid $a$ over a time window of 5000 MCMC steps after $t_{\rm MUT}$. The faded area is the standard deviation.